OpenStudy (anonymous):

A school has 2000 students. 1200 students are girls what is the ratio of ratio of boys to girls?

OpenStudy (anonymous):

2000-1200=800 boys so 800/1200=2/3 so 2:3

OpenStudy (saifoo.khan):

Roshan got that.

OpenStudy (anonymous):

thank you. What about this one. The rams won 8 games and lost 6 games last season. What is the ratio of wins and loses. would it be 6:8?

OpenStudy (saifoo.khan):

3:4

OpenStudy (anonymous):

Can you work it out for me

OpenStudy (saifoo.khan):

k

OpenStudy (saifoo.khan):

6 : 8 Divide both by 2 3 : 4

OpenStudy (anonymous):

Gotcha thanks.

OpenStudy (anonymous):

(2000-1200) : 1200 = 800 :1200 =8 :12 = 2:3